Kitchen Supply, Inc. (KSI), manufactures three types of flatware: institutional, standard, and silver. It applies all indirect costs according to a predetermined rate based on direct labor-hours. A consultant recently suggested that the company switch to an activity-based costing system and prepared the following cost estimates for year 2 for the recommended cost drivers.

Activity

Recommended

Cost Diver

Estimated

Cost

Estimated Cost

Driver Activity

Processing orders Number of orders $ 46,000 200 orders
Setting up production Number of production runs 180,000 100 runs
Handling materials Pounds of Materials used 220,000 110,000 pounds

Machine depreciation and

maintenance

Machine-hours 220,000 11,000 hours
Performing quality control Number of inspections 62,100 45 inspections
Packing Number of units 135,000 450,000 units
Total estimated cost $863,100

In addition, management estimated 7,300 direct labor-hours for year 2.

Assume that the following cost driver volumes occurred in January, year 2.

Institutional Standard Silver
Number of units produced 60,000 21,000 8,000
Direct materials costs $38,000 $20,000 $15,000
Direct labor-hours 400 410 630
Number of orders 13 7 6
Number of production runs 3 3 6
Pounds of material 14,000 6,000 3,200
Machine-hours 560 140 80
Number of inspections 4 3 2
Units shipped 60,000 21,000 8,000

Actual labor costs were $16 per hour.

Required: a. (1) Compute a predetermined overhead rate for year 2 for each cost driver using the estimated costs and estimated cost driver units prepared by the consultant.

(2) Compute a predetermined rate for year 2 using direct labor-hours as the allocation base.

b. Compute the production costs for each product for January using direct labor-hours as the allocation base and the predetermined rate computed in requirement a(2).

c. Compute the production costs for each product for January using the cost drivers recommended by the consultant and the predetermined rates computed in requirement a. (Note: Do not assume that total overhead applied to products in January will be the same for activity-based costing as it was for the labor-hour-based allocation.)
Answer #1

Answer with working notes is given below

Calculation of Activity rates Total Total Activity activity rate Processing order Setting up production Handling material Mac
Answer #2

B)institutional

400*118.23= 47292

400*16= 6400

indirect cost (total overhead allocated)


230*13= 2,990